OpenStudy (coolguy34):

i need help, please give me the step by step solution to 2(x – 3) – 1 = -10

OpenStudy (anonymous):

2x-6-1=-10 2x-7=-10 |+7 2x=-3 X=-3/2

OpenStudy (anonymous):

1. Add 1 to both sides: 2(x-3) = -9 2. Multiply left side out: 2x -6 = -9 3. Add 6 to both sides: 2x = -3 4. Divide by 2 both sides: x = -3/6 5. Simplify: x = -1/2

OpenStudy (coolguy34):

thanks @someone7070